Planning Notebook

by Kenneth John Odle

v. 0.1.0 3 July 2021 2

Contents

1 Introduction 1 1.1 Origins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 An Important Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 A Minor Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Goals 3 3 Annual Goals 5 4 Weekly Tasks 7 5 The Year 9 4

Chapter 1

Introduction

1.1 Origins

I created this for one main reason: I have a hard time getting and staying organized, especially with regard to financial matters. Part of this is no doubt due to my background (I did not grow up in an environment where people moved money around to benefit themselves) and undiagnosed ADHD (which is something I’m also looking into).

Also, I had a hard time making this and making it look nice (I normally do a lot of design work when I have the opportunity) with the tools I had available to me. Because I like learning new things, I decided to finally learn how to use L^{A}TEX. (I have no idea what took me so long—the learning curve is not steep and it’s remarkably powerful.) Because it’s basically a text file, I can compile it to a pdf file to share, but I can also share it via Git for those who want to create their own version.

1.2 An Important Note

You don’t have to wait until the beginning of a new year to start this. In fact, I suggest that you avoid New Year’s Day to make any major changes in your life. There is already so much going on, what with it being the height of the holiday season for most western religions. As well, people expect you to commit to changes. (And take delight in when you fail.)

Rather, start where you are, and make changes as gradually or as quickly as you feel comfortable with.

1.3 A Minor Note

Like I mentioned above, I’ve generated this document using L^{A}TEX. As a result, this document is subject to all the limitations of L^{A}TEX, and more specifically, my limited (though growing) abilities with L^{A}TEX. If you want to correct any markup errors I’ve made (and I’m sure there are plenty), please contact me.

2

Chapter 2

Goals

Let’s talk about goals.

Presumably you want to make a plan because you have some end in mind—something in your life that you want to change, see more of, or see less of. We’re going to work on a year-long plan to get you at least partway to those goals.

4

Chapter 3

Annual Goals

Okay, let’s make some goals for the next 365 days.

6

Chapter 4 Weekly Tasks

7 8

Chapter 5

The Year

The rest of this planner consists of 52 weeks of planning pages.

Month and Year Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su Goals for this Year Goals for this Month